Q36: Megan purchases a perpetuity-immediate for 3250 with annual payments of 130. At the same price...

Q36:

Megan purchases a perpetuity-immediate for 3250 with annual payments of 130. At the same price and interest rate, Chris purchases an annuity-immediate with 20 annual payments that begin at amount P and increase by 15 (growth) each year thereafter. Calculate P .

Correct answer is: 116
Answer #1

Present value of perpetuity = Annual payment / Interest rate
3250 = 130 / Interest rate
Interest rate = 130 / 3250 = 0.04 or 4%

r = 0.04
n = 20
D = 15

(1/r) * (1 - (1 / (1 + r)^n)) = (1/0.04) * (1 - (1/(1+0.04)^20))
= 25 * (1 - 0.4564)
= 13.5903

PV = 3250 = (P * 13.5903) + (15 * (13.5903 - 20/(1 + 0.04)^20) / 0.04
3250 = (P * 13.5903) + (15 * (13.5903 - 9.1277) / 0.04
3250 = (P * 13.5903) + (15 * (4.4626 / 0.04)
3250 = (P * 13.5903) + 1673.4750
P = (3250 - 1673.4750) / 13.5903
P = 1576.5250 / 13.5903
P = 116